What is the boiling point of a solution that is produced by dissolving 20g NaOH in 2L H2O if the Kb of H2O is 0.52 oC/m?

a. 100.0oC
b. 100.3 oC
c. 100.4 oC
d. 101.0 oC
1
Expert's answer
2020-04-23T13:09:16-0400

"\Delta T=K_b *m_l *i" , "\Delta T" - temperature change (T2-T1), Kb - the molal boiling point elevation constant, ml - molality or molal concentration of the solute in the solution, i - isotonic coefficient that takes into account dissociation; "i = 1+\alpha (\nu -1)", "\alpha" - degree of dissociation, "\nu" - the number of ions formed during dissociation. "n=m\/M" , n - chemical amount of substance; m - mass of substance; M - molar mass of substance. M(NaOH)=23g/mol+16g/mol+1g/mol=40g/mol.n(NaOH)=20g/40g/mol=0,5mol"m_l =n\/m(H_2O)" , n - chemical amount of substance; m(H2O) - solvent mass(kg), in our case - H2O. "m=V*\rho" , m - mass, V - volume, "\rho" - density (for H2O is 997kg/m3).2L=2dm3=0,002m3 m(H2O)=0,002m3*997kg/m3=1,994kgml=0,5mol/1,994kg=0,250mol/kg"i" = 1+1(2-1)=2 ("NaOH\to Na^+ +OH^-" , "\nu" =2, "\alpha" =1, because NaOH is a strong electrolyte)

"\Delta T = 0,52oC/m*0,250mol/kg*2=0,26"oC

T1=100oC (water boiling point), "T_2 =T_1 +\Delta T" , T2=100,26 round the number to 100,3.

Answer: b.